# Exact Sequence Interpolation with Transformers

### v2

@@
Abstract.
We prove that transformers can exactly interpolate datasets of finite input sequences in $\mathbb{R}^{d}$, $d\geq 2$, with corresponding output sequences of smaller or equal length. Specifically, given $N$ sequences of arbitrary but finite lengths in $\mathbb{R}^{d}$ and output sequences of lengths $m^{1},\dots,m^{N}\in\mathbb{N}$, we construct a transformer with $\mathcal{O}(\sum_{j=1}^{N}m^{j})$ blocks and $\mathcal{O}(d\sum_{j=1}^{N}m^{j})$ parameters that exactly interpolates the dataset. Our construction provides complexity estimates that are independent of the input sequence length, by alternating feed-forward and self-attention layers and by capitalizing on the clustering effect inherent to the latter. Our novel constructive method also uses low-rank parameter matrices in the self-attention mechanism, a common feature of practical transformer implementations. These results are first established in the hardmax self-attention setting, where the geometric structure permits an explicit and quantitative analysis, and are then extended to the softmax setting. Finally, we demonstrate the applicability of our exact interpolation construction to learning problems, in particular by providing convergence guarantees to a global minimizer under regularized training strategies. Our analysis contributes to the theoretical understanding of transformer models, offering an explanation for their excellent performance in exact sequence-to-sequence interpolation tasks.
Chair for Dynamics, Control, Machine Learning, and Numerics (Alexander von Humboldt Professorship), Department of Mathematics, Friedrich–Alexander–Universität Erlangen–Nürnberg, 91058 Erlangen, Germany.
*Departamento de Matemáticas, Universidad Autónoma de Madrid, 28049 Madrid, Spain.
Chair of Computational Mathematics, Fundación Deusto. Av. de las Universidades, 24, 48007 Bilbao, Basque Country, Spain.
1. Introduction
Transformers [31] have revolutionized machine learning (ML) by outperforming traditional residual networks (ResNets) in applications such as natural language processing [25] and computer vision [8], where inputs naturally take the form of finite sequences of $d$-dimensional vectors. Their practical success relies on the effectiveness of the self-attention layers, which act between feed-forward layers to update each vector in the sequence by aggregating information from all other vectors.
The expressive power of feed-forward layers has already been studied extensively, see, e.g., [6, 17, 21, 27, 32, 26, 30, 22, 29, 35]. In particular, it is known [33] that sufficiently deep feed-forward networks can interpolate arbitrary input data. Unlike feed-forward networks, transformers rely on self-attention, and thus their interpolation properties and complexity scaling require a distinct analysis.
1.1. Summary of Contributions
@@
  • Exact interpolation with real vector-valued outputs. We show via an explicit construction that transformer parameters can always be chosen to interpolate arbitrary real vector-valued sequence-to-sequence maps exactly. We therefore overcome previous limitations to discrete outputs [19] or approximate interpolation [12] (see also Section˜1.4 below for further discussion).
@@
  • Parameter efficiency through low-rank attention. By parametrizing the attention layer of our transformer using low-rank matrices or multiples of the identity, the number of parameters in our construction grows linearly with the input embedding dimension $d$. This is consistent with efficient transformer designs used in practice.
@@
We present the problem statement and our results more precisely in Section˜1.2 and Section˜1.3. Similarities and differences with existing works are discussed in Section˜1.4. We conclude by comparing the parameter complexity of transformers and ResNets in Section˜1.5.
1.2. Problem Statement
Fix $d\in\mathbb{N}$. For $n\in\mathbb{N}$, we denote by $(\mathbb{R}^{d})^{n}$ the set of tuples of vectors in $\mathbb{R}^{d}$. In keeping with standard terminology, we call $X=(x_{1},\ldots,x_{n})\in(\mathbb{R}^{d})^{n}$ a sequence and each vector $x_{i}$ a token. Since our constructions are permutation-equivariant [34], we identify two sequences if they differ only by a permutation of their tokens, so sequences may be regarded mathematically as unordered sets. The set of all finite-length non-empty sequences in $\mathbb{R}^{d}$ is
@@ 1.2. Problem Statement
(1.1) $d_{H}(X,Y)\coloneqq\max\left(\max_{x\in X}\min_{y\in Y}\|x-y\|_{2},\max_{y\in Y}\min_{x\in X}\|x-y\|_{2}\right),$
@@
Sequential datasets with the previously described characteristics appear naturally in language and vision applications. In masked language modeling [7], each input sequence $X^{j}$ corresponds to an incomplete sentence and the output $Y^{j}$ consists of the missing (or “masked”) words. Analogously, in masked image modeling [14, 4], the input sequence $X^{j}$ represents an image with missing patches, while the output $Y^{j}$ contains the missing content. For further illustration, we consider the following example in language modeling.
@@
The interpolation problem for the dataset consists of constructing a function that maps each input $X^{j}$ to the corresponding output $Y^{j}$. Transformers are an appropriate class of models to solve such problems, since they process variable-length inputs in a permutation-equivariant way. For our purposes, we view a transformer as a parametrized map $\operatorname{\mathrm{T}}:\mathcal{X}\to\mathcal{X}$ acting on sequences of vectors in $\mathbb{R}^{d}$. Its precise architecture will be described in Section˜3, but for the present discussion it suffices to note that transformers preserve sequence length: if $X^{j}$ has length $n^{j}$, then $\operatorname{\mathrm{T}}(X^{j})$ also has length $n^{j}$. This creates a mismatch with our outputs $Y^{j}$, which are of equal or shorter length ($m^{j}\leq n^{j}$). We resolve this mismatch by viewing equality between sequences in the set sense, that is, two sequences are equal when their Hausdorff distance is equal to zero. We are therefore led to the following exact interpolation problem.
@@
To ensure this problem is solvable, we assume the following on the dataset.
@@
The input and output sequences $\{(X^{j},Y^{j})\}_{j\in[N]}$ satisfy that
@@
    the sequences $X^{1},\ldots,X^{N}$ are pairwise distinct,
@@
    tokens within each input and output sequence are pairwise distinct.
Assumption i) is very mild: if $X^{j}=X^{j^{\prime}}$ and $Y^{j}=Y^{j^{\prime}}$, then one of the sequences can be dropped without changing ˜1.2. If instead, $X^{j}=X^{j^{\prime}}$ and $Y^{j}\neq Y^{j^{\prime}}$, then ˜1.2 trivially has no solution. Assumption ii) is also mild since positional encoding [28] usually ensures that tokens in a sequence are distinct. The pairwise-distinctness requirement is not essential: with minor proof edits (at the cost of heavier notation), repeated tokens in an input sequence $X^{j}$ can be replaced by a single token with more “mass”. Concretely, replace the constant $1/|\mathcal{C}_{i}(X,A)|$ in ˜3.3 with $w_{\ell}/\sum_{k\in\mathcal{C}_{i}(X,A)}w_{k}$, where $w_{\ell}$ counts the number of times token $x_{\ell}$ is repeated.
1.3. Main Results
We solve ˜1.2 for datasets satisfying ˜1.3 in dimension $d\geq 2$ [^1] The case $d=1$ is of limited practical interest, so the restriction is mild. using a transformer that alternates single-head self-attention and feed-forward layers with ReLU activations and trainable residual connections. We refer to a single pair of feed-forward and self-attention layers as a transformer block. Details of the architecture are described in Section˜3.
We analyze two cases, depending on whether the self-attention layers are defined using a hardmax or a softmax attention mechanism. We begin with the hardmax case, which, although not used in practice, offers a simpler geometric structure that facilitates the analysis [2]. Then, we extend the result to the softmax case, which is in line with standard transformer architectures. The two cases are closely related, since the hardmax function arises as a temperature parameter $\tau>0$ in the softmax function tends to zero [9].
@@ 1.3. Main Results
Fix $N\in\mathbb{N}$, $d\geq 2$ and a dataset of sequences $\{(X^{j},Y^{j})\}_{j\in[N]}$ satisfying ˜1.3. There exists a hardmax transformer $\operatorname{\mathrm{T}}^{0}$ with feed-forward layers of width $d^{\prime}\leq 4$,
(1.2) $L=2\sum_{j=1}^{N}m^{j}+2N+1\;\text{blocks,}\quad\text{and}\quad P=\mathcal{O}\left(d\sum_{j=1}^{N}m^{j}\right)\;\text{parameters}$
@@
Our hardmax transformer has $\mathcal{O}(d\sum_{j=1}^{N}m^{j})$ parameters, rather than $\mathcal{O}(d^{2}\sum_{j=1}^{N}m^{j})$ that one might expect at first, because in every self-attention layer we choose the tunable matrix parameters either as scalar multiples of the identity, or as rank-one matrices of the form $vv^{\top}$ for some $v\in\mathbb{R}^{d}$. In addition, the parameters in the feed-forward layers are always described by vectors in $\mathbb{R}^{d}$ and $d\times d^{\prime}$ matrices for $d^{\prime}\leq 4$.
The proof of Theorem˜1.4 is presented in Section˜5. In Section˜6, instead, we prove that the result remains true for the softmax case at the expense of a slight increase in the number of blocks.
@@
Fix $N\in\mathbb{N}$, $d\geq 2$ and a dataset of sequences $\{(X^{j},Y^{j})\}_{j\in[N]}$ satisfying ˜1.3. There exists $\tau^{\star}>0$ such that, for all $0<\tau<\tau^{\star}$, there exists a softmax transformer $\operatorname{\mathrm{T}}^{\tau}$ with feed-forward layers of width $d^{\prime}\leq 4$,
(1.3) $L=2\sum_{j=1}^{N}m^{j}+3N\;\text{blocks,}\quad\text{and}\quad P=\mathcal{O}\left(d\sum_{j=1}^{N}m^{j}\right)\;\text{parameters}$
@@
Several comments about Theorem˜1.6 are in order. First, to the best of our knowledge this is the first exact interpolation result for softmax transformers with real-valued outputs (see Section˜1.4 for further discussion). Moreover, we derive explicit complexity bounds in terms of both the number of blocks and the total parameter count.
@@
Second, the theorem does not follow directly from the hardmax result by taking the “zero-temperature limit” of the softmax [9]. Such an argument would only yield approximate interpolation, with errors depending on $\tau>0$. Our key contribution is to eliminate these errors exactly by exploiting the properties of ReLU feed-forward layers, at the cost of only N1N-1 additional blocks.
@@
Finally, the temperature can be normalized to $\tau^{\star}=1$ by rescaling the self-attention parameter matrix in (3.2). However, this rescaling comes at the expense of increasing the norm of the parameters in the self-attention layer. Bounding parameter norms is in fact substantially more delicate, and our results do not attempt to do so; in practice, such norms depend nontrivially on the dispersion of the data and the separation of the outputs.
1.4. Related Work
@@
The interpolation capacity of transformers has been studied in previous works. In [19], exact interpolation results were established for sequence-to-sequence maps with equal input and output lengths ($m=n$) and integer-valued outputs. The estimates in that work show that the required number of blocks grows as $\tilde{\mathcal{O}}(n+\sqrt{Nn})$, where the $\tilde{\mathcal{O}}$ notation hides logarithmic factors in $N$, $n$, and $d$. For real-valued outputs, however, the analysis yields only approximate interpolation [19, Remark 3.2]. In contrast, our results prove exact interpolation for real-valued vector outputs. This requires a number of blocks that grows linearly in both $N$ and $n$, but crucially remains independent of $d$.
@@
A different line of work is pursued in [23], which studies interpolation in transformers with a single multi-head attention layer. There, the number of heads plays a role analogous to the width of feed-forward layers, and the authors show that $N$ sequences of length $n$ can be memorized with at least $\mathcal{O}(N/n)$ heads. Our analysis focuses instead on single-head deep transformers, a structurally different architecture that leads to complementary results (see Section˜7.2.3 for further discussion on the multi-head case).
@@
Beyond interpolation, several works have investigated the function approximation capabilities of transformers. The seminal result of [34] established that transformers are universal approximators, while [1] extended these results to efficient variants of transformer architectures. A more recent perspective is offered by [12], where transformers are interpreted as interacting particle systems within a measure-theoretic framework. In this setting, transformers achieve approximate interpolation in the Wasserstein-2 metric for measure-to-measure maps. The construction in [12] can also solve ˜1.2, but requires $\mathcal{O}(Nd+\sum_{j=1}^{N}m^{j})$ blocks and is limited to the case of datasets where inputs and outputs have the same length $m^{j}=\operatorname{len}(X^{j})=\operatorname{len}(Y^{j})$. Theorem˜1.6 improves upon these works by achieving lower block complexity and accommodating arbitrary output sequence lengths.
1.5. Parameter Complexity of Transformers versus ResNets

### v3

@@
Abstract
We prove that transformers can exactly interpolate datasets of finite input sequences in $\mathbb{R}^{d}$, $d\geq 2$, with corresponding output sequences of smaller or equal length. Specifically, given $N$ sequences of arbitrary but finite lengths in $\mathbb{R}^{d}$ and output sequences of lengths $m^{1},\dots,m^{N}\in\mathbb{N}$, we construct a transformer with $\mathcal{O}(\sum_{j=1}^{N}m^{j})$ blocks and $\smash{\mathcal{O}(d\sum_{j=1}^{N}m^{j})}$ parameters that exactly interpolates the dataset. Our construction provides complexity estimates that are independent of the input sequence length, by alternating feed-forward and self-attention layers and by capitalizing on the clustering effect inherent to the latter. Our novel constructive method also uses low-rank parameter matrices in the self-attention mechanism, a common feature of practical transformer implementations. These results are first established in the hardmax self-attention setting, where the geometric structure permits an explicit and quantitative analysis, and are then extended to the softmax setting. Finally, we demonstrate the applicability of our exact interpolation construction to learning problems, in particular by providing convergence guarantees to a global minimizer under regularized training strategies. Our analysis contributes to the theoretical understanding of transformer models, offering an explanation for their excellent performance in exact sequence-to-sequence interpolation tasks.
keywords:
Deep Learning, Self-Attention, Approximation, Transformer Architecture, Mathematics of Machine Learning
footnotetext: Author accepted manuscript. The version of record is published in Math. Found. Mach. Learn. 2, 2 (2026), doi:10.1007/s44439-026-00005-y.
1 Introduction
Transformers [vaswaniAttentionAllYou2017] have revolutionized machine learning (ML) by outperforming traditional residual networks (ResNets) in applications such as natural language processing [openai2024gpt4technicalreport] and computer vision [dosovitskiy2021imageworth16x16words], where inputs naturally take the form of finite sequences of $d$-dimensional vectors. Their practical success relies on the effectiveness of the self-attention layers, which act between feed-forward layers to update each vector in the sequence by aggregating information from all other vectors.
The expressive power of feed-forward layers has already been studied extensively, see, e.g., [Cybenko1989ApproximationBS, HORNIK1991251, leshno1993multilayer, Pinkus_1999, yarotsky2017error, park2021provable, vardi2022on, li2022deep, domenec2023NODES, alvarez2025clusterClassification]. In particular, it is known [yun2019small] that sufficiently deep feed-forward networks can interpolate arbitrary input data. Unlike feed-forward networks, transformers rely on self-attention, and thus their interpolation properties and complexity scaling require a distinct analysis.
1.1 Summary of contributions
@@
  • Exact interpolation with real vector-valued outputs. We show via an explicit construction that transformer parameters can always be chosen to interpolate arbitrary real vector-valued sequence-to-sequence maps exactly. We therefore overcome previous limitations to discrete outputs [kim2023provable] or approximate interpolation [geshkovski2024measure] (see also Section 1.4 below for further discussion).
@@
  • Parameter efficiency through low-rank attention. By parameterizing the attention layer of our transformer using matrices that are sums of a rank-one matrix and a multiple of the identity, the number of parameters in our construction grows linearly with the input embedding dimension $d$. This is consistent with efficient transformer designs used in practice.
@@
We present the problem statement and our results more precisely in Section 1.2 and Section 1.3. Similarities and differences with existing works are discussed in Section 1.4. We conclude by comparing the parameter complexity of transformers and ResNets in Section 1.5.
1.2 Problem statement
Fix $d\in\mathbb{N}$. For $n\in\mathbb{N}$, we denote by $(\mathbb{R}^{d})^{n}$ the set of tuples of vectors in $\mathbb{R}^{d}$. In keeping with standard terminology, we call $X=(x_{1},\ldots,x_{n})\in(\mathbb{R}^{d})^{n}$ a sequence and each vector $x_{i}$ a token. Since our constructions are permutation-equivariant [yunAreTransformersUniversal2020], we identify two sequences if they differ only by a permutation of their tokens, so sequences may be regarded mathematically as unordered sets. The set of all finite-length non-empty sequences in $\mathbb{R}^{d}$ is
@@ 1.2 Problem statement
$d_{H}(X,Y)\coloneqq\max\left(\max_{x\in X}\min_{y\in Y}\|x-y\|_{2},\max_{y\in Y}\min_{x\in X}\|x-y\|_{2}\right),$ (1.1)
@@
Sequential datasets with the previously described characteristics appear naturally in language and vision applications. In masked language modeling [devlin-etal-2019-bert], each input sequence $X^{j}$ corresponds to an incomplete sentence and the output $Y^{j}$ consists of the missing (or “masked”) words. Analogously, in masked image modeling [he2022masked, bao2022beit], the input sequence $X^{j}$ represents an image with missing patches, while the output $Y^{j}$ contains the missing content. For further illustration, we consider the following example in language modeling.
@@
The interpolation problem for the dataset consists of constructing a function that maps each input $X^{j}$ to the corresponding output $Y^{j}$. Transformers are an appropriate class of models to solve such problems, since they process variable-length inputs in a permutation-equivariant way. For our purposes, we view a transformer as a parametrized map $\operatorname{\mathrm{T}}:\mathcal{X}\to\mathcal{X}$ acting on sequences of vectors in $\mathbb{R}^{d}$. Its precise architecture will be described in Section 3, but for the present discussion it suffices to note that transformers preserve sequence length: if $X^{j}$ has length $n^{j}$, then $\operatorname{\mathrm{T}}(X^{j})$ also has length $n^{j}$. This creates a mismatch with our outputs $Y^{j}$, which are of equal or shorter length ($m^{j}\leq n^{j}$). We resolve this mismatch by viewing equality between sequences in the set sense, that is, two sequences are equal when their Hausdorff distance is equal to zero. We are therefore led to the following exact interpolation problem.
@@
Remark 1.3.
1.2 is formulated for permutation-equivariant transformers, where the order of tokens does not matter. Positional encodings [vaswaniAttentionAllYou2017] can be employed in a separate pre-processing step to add information about the order of tokens if desired, but we do not account for it here because this does not affect the behavior of the transformer. For sequence modeling tasks that require auto-regressive next-token prediction, the permutation equivariance can also be removed by employing “masked” self-attention. The relation between this masked auto-regressive setting and 1.2 is discussed in Section 7.2.4, where we show how the former can be reduced to an instance of the latter.
Remark 1.4.
Viewing the equality $\operatorname{\mathrm{T}}(X^{j})=Y^{j}$ as a set equality means that, if the output sequence consists of a single token, then every token of the sequence $\operatorname{\mathrm{T}}(X^{j})$ must coincide with the desired output token. This is harder to achieve compared to matching only a specific token of the input sequence to the desired output, but can be done with a transformer with the same leading-order complexity.
To ensure 1.2 is solvable, we make the following assumptions on the dataset.
@@
The input and output sequences $\{(X^{j},Y^{j})\}_{j\in[N]}$ have the following properties:
@@
    The sequences $X^{1},\ldots,X^{N}$ are pairwise distinct.
@@
    Tokens within each $X^{j}$ are pairwise distinct.
  3. iii)
    Tokens within each $Y^{j}$ are pairwise distinct.
Assumption i) could be replaced with the weaker requirement that $X^{j}=X^{j^{\prime}}$ implies $Y^{j}=Y^{j^{\prime}}$. This condition is necessary to ensure that 1.2 is solvable. However, if $X^{j}=X^{j^{\prime}}$ and $Y^{j}=Y^{j^{\prime}}$, then the pair $(X^{j^{\prime}},Y^{j^{\prime}})$ can be dropped from the dataset without changing 1.2 to ensure that Assumption i) holds. Assumptions ii) and iii) are mild since positional encoding [press2022trainshorttestlong] usually ensures that tokens in a sequence are distinct. The pairwise-distinctness requirement is not essential: with minor proof edits (at the cost of heavier notation), repeated tokens in an input sequence $X^{j}$ can be replaced by a single token with more “mass”. Concretely, replace the constant $1/|\mathcal{C}_{i}(X,A)|$ in 3.3 with $w_{\ell}/\sum_{k\in\mathcal{C}_{i}(X,A)}w_{k}$, where $w_{\ell}$ counts the number of times token $x_{\ell}$ is repeated.
1.3 Main results
We solve 1.2 for datasets satisfying 1.5 in dimension $d\geq 2$ [^0] The case $d=1$ is of limited practical interest, so the restriction is mild. using a transformer with $\smash{\mathcal{O}(d\sum_{j}m_{j})}$ parameters that alternates single-head self-attention and feed-forward layers with ReLU activations and trainable residual connections. We refer to a single pair of feed-forward and self-attention layers as a transformer block. Details of the architecture are described in Section 3.
We analyze two cases, depending on whether the self-attention layers are defined using a hardmax or a softmax attention mechanism. While only the latter is used in practice, the former allows for a simpler geometric intuition that facilitates analysis [alcalde2025clustering]. Additionally, “hardmax transformers” are closely related to “softmax transformers” because the hardmax function arises when the so-called temperature parameter in the softmax function tends to zero [elfadelSoftmax1993].
Our first main result establishes that hardmax transformers can exactly interpolate data and solve 1.2.
@@ 1.3 Main results
Fix $N\in\mathbb{N}$, $d\geq 2$ and a dataset of sequences $\{(X^{j},Y^{j})\}_{j\in[N]}$ satisfying 1.5. There exists a hardmax transformer $\operatorname{\mathrm{T}}^{0}$ with feed-forward layers of width $d^{\prime}\leq 4$,
$L=2\sum_{j=1}^{N}m^{j}+2N+1\;\text{blocks,}\quad\text{and}\quad P=\mathcal{O}\left(d\sum_{j=1}^{N}m^{j}\right)\;\text{parameters}$ (1.2)
@@
The number of parameters in Theorem 1.6 agrees up to a prefactor with the lower bound of $\smash{P\geq d\sum_{j}m^{j}}$ for any function $\operatorname{\mathrm{T}}(\cdot,\theta):\mathcal{X}\to\mathcal{X}$ parameterized in a smooth way by $\theta\in\mathbb{R}^{P}$. Indeed, if a parameter vector $\theta$ exists for every choice of output sequences $\{Y^{j}\}$ in a $\smash{d\sum_{j}m^{j}}$-dimensional subspace $\mathcal{Y}$ of finite-length sequences, then the map from the parameter space $\mathbb{R}^{P}$ to this space must be surjective. This requires $\smash{P\geq d\sum_{j}m^{j}}$, otherwise every point in $\mathbb{R}^{P}$ would be critical and the image $\operatorname{\mathrm{T}}(\cdot,\mathbb{R}^{P})\subset\mathcal{Y}$ would have measure zero by Sard’s theorem [Guillemin1974, §7].
The proof of Theorem 1.6, presented in Section 5, relies on a number of technical results proven in Section 4 that may be of independent interest. In particular, Lemma 4.4 shows how a self-attention layer can collapse tokens that lie in a suitable geometric configuration. Conversely, Lemmas 4.7 and 4.8 show how self-attention and feed-forward layers can be used to separate overlapping tokens from different sequences. Finally, Lemma 4.8 demonstrates that feed-forward layers can move tokens individually once they are separated and collapsed. Together, these results provide simple geometric tools that underlie the construction of the interpolating transformer.
The second main result of this paper, proven in Section 6, is a solution to the exact interpolation problem that uses softmax transformers. Specifically, we prove that the statement of Theorem 1.6 remains true for softmax transformers at the expense of a slight increase in the number of blocks.
@@
Fix $N\in\mathbb{N}$, $d\geq 2$ and a dataset of sequences $\{(X^{j},Y^{j})\}_{j\in[N]}$ satisfying 1.5. There exists $\tau^{\star}>0$ such that, for all $0<\tau<\tau^{\star}$, there exists a softmax transformer $\operatorname{\mathrm{T}}^{\tau}$ with feed-forward layers of width $d^{\prime}\leq 4$,
$L=2\sum_{j=1}^{N}m^{j}+3N\;\text{blocks,}\quad\text{and}\quad P=\mathcal{O}\left(d\sum_{j=1}^{N}m^{j}\right)\;\text{parameters}$ (1.3)
@@
Several comments about Theorem 1.8 are in order. First, to the best of our knowledge this is the first exact interpolation result for deep softmax transformers with real-valued outputs (see Section 1.4 for further discussion). Moreover, we derive explicit complexity bounds in terms of both the number of blocks and the parameter count.
@@
Second, the theorem does not follow directly from the hardmax result by taking the “zero-temperature limit” of the softmax [elfadelSoftmax1993]. Such an argument would only yield approximate interpolation, with errors depending on $\tau>0$. Our key contribution is to eliminate these errors exactly by exploiting the properties of ReLU feed-forward layers, at the cost of only N1N-1 additional blocks.
@@
Finally, the temperature can be normalized to $\tau^{\star}=1$ by rescaling the self-attention parameter matrix in (3.2). However, this rescaling comes at the expense of increasing the norm of the parameters in the self-attention layer. Bounding parameter norms is a substantially more delicate problem, which we do not attempt to solve. In practice, such norms depend nontrivially on the dispersion of the data and the separation of the outputs [kim2023provable]. See Section 7.2.1 for further discussion.
1.4 Related work
@@
The interpolation capacity of transformers has been studied in previous works. In [kim2023provable], exact interpolation results were established for sequence-to-sequence maps with equal input and output lengths ($m=n$) and integer-valued outputs. The estimates in that work show that the required number of blocks grows as $\tilde{\mathcal{O}}(n+\sqrt{Nn})$, where the $\tilde{\mathcal{O}}$ notation hides logarithmic factors in $N$, $n$, and $d$. For real-valued outputs, however, the analysis yields only approximate interpolation [kim2023provable, Remark 3.2]. In [kajitsuka2024are], the analysis of [kim2023provable] is extended to vector-valued discrete output sequences using one-layer self-attention with low-rank weight matrices. Regarding next-token prediction capacity of shallow decoder transformers, [madden2025next] prove an exact interpolation result where outputs are sequences of probability vectors over a finite vocabulary (see Section 7.2.2 for further comments on this setting). The interpolation properties of transformers have also been analyzed from a controllability-based viewpoint in [agrachev2025generic], where it is proven that exact interpolation is possible using transformers with more general self-attention layers than those commonly used in practice. In contrast, we prove exact interpolation for real-valued vector outputs in sequence-to-sequence tasks using deep single-head transformers with the standard self-attention mechanism. This requires a number of blocks that grows linearly in both $N$ and $n$, but crucially remains independent of $d$.
@@
A different line of work is pursued in [mahdavi2024memorization], which studies interpolation in transformers with a single multi-head attention layer. There, the number of heads plays a role analogous to the width of feed-forward layers, and the authors show that $N$ sequences of length $n$ can be memorized with at least $\mathcal{O}(N/n)$ heads. Our analysis focuses instead on single-head deep transformers, a structurally different architecture that leads to complementary results (see Section 7.2.3 for further discussion on the multi-head case).
@@
Finally, beyond interpolation, there are several works investigating the function approximation capabilities and dynamical behavior of transformers. For instance, the asymptotic dynamics of hardmax attention have been shown to induce clustering phenomena in token representations [alcalde2025clustering]. The seminal result of [yunAreTransformersUniversal2020] established that transformers are universal approximators of sequence-to-sequence maps. The work [alberti2023sumformer] extends these results to efficient variants of transformers, while [furuya2025transformers] proves that transformers are universal approximators of “in-context maps” (see that work for a definition of this term). Yet another perspective is offered by [geshkovski2024measure], where transformers are interpreted as interacting particle systems within a measure-theoretic framework. In this setting, transformers achieve approximate interpolation in the Wasserstein-2 metric for measure-to-measure maps. The construction in [geshkovski2024measure] can also solve 1.2, but requires $\mathcal{O}(Nd+\sum_{j=1}^{N}m^{j})$ blocks and is limited to the case of datasets where inputs and outputs have the same length $m^{j}=\operatorname{len}(X^{j})=\operatorname{len}(Y^{j})$. Theorem 1.8 improves upon these works by achieving lower block complexity and accommodating arbitrary output sequence lengths.
1.5 Parameter complexity of transformers versus ResNets